# Decidability of polynomial equations over function fields in positive characteristic

### v1

@@ 1. Introduction
This question, too, remains unresolved to this day. Anscombe and Fehm recently established a positive answer to a variant of 1.1, with $K(t)$ replaced by the field of formal Laurent series K((t))K(\!(t)\!) [AF16].
@@
Here, one may remark that undecidability of $\operatorname{Th}_{\exists}(\overline{\mathbb{F}_{p}}(t_{1},t_{2}),\mathbb{Z}[t_{1},t_{2}])$ was shown by Kim and Roush for $p>2$ [KR92], later generalized to include finite extensions of $\overline{\mathbb{F}_{p}}(t_{1},t_{2})$ as well by Eisenträger [Eis12]. Recently Becher, Dittmann and the author obtained more examples of function fields $F/K$ where $K$ contains $\overline{\mathbb{F}_{p}}$ and $\operatorname{Th}_{\exists}(F,R)$ is shown to be undecidable for some well-chosen ring of coefficients $R$ [BDD23], relying partially on the type of reduction techniques discussed in the present paper.
@@ 2. Preliminaries on definability
Consider algebraic field extensions $k\subseteq l\subseteq m$ such that $l/k$ is Galois, $m/l$ is finite simple, and $m/k$ is normal. Further consider an arbitrary field extension $L/l$ which is linearly disjoint from $m$ over $l$, and denote $M=L\otimes_{l}m$. If $S\subseteq M^{n}$ is positive-existentially ${\mathcal{L}_{\rm ring}}(k)$-definable over $M$, then $S\cap L^{n}$ is positive-existentially ${\mathcal{L}_{\rm ring}}(k)$-definable over $L$.
@@
Let $\alpha\in m$ be a generator of $m$ over $l$. Let $Q_{1}(X)\in l[X]$ be the minimal polynomial of $\alpha$ over $l$, and denote by $Q_{2}(X),\ldots,Q_{g}(X)$ the different $k$-conjugates of $Q_{1}(X)$. Since $L$ is linearly disjoint from $m$ over $l$, each $Q_{i}(X)$ is irreducible over $L$. Let $P(X)=\prod_{i=1}^{g}Q_{i}(X)$ – as $l/k$ is Galois, we have that $P(X)$ is the minimal polynomial of $\alpha$ over $k$. Define the $L$-algebra $\mathcal{D}=\frac{L[X]}{(P(X))}$.
Now consider a positive-existential ${\mathcal{L}_{\rm ring}}(k)$-formula $\varphi(x_{1},\ldots,x_{n})$ such that $S=\{(b_{1},\ldots,b_{n})\in M^{n}\mid M\models\varphi(b_{1},\ldots,b_{n})\}$. We claim that
$S\cap L^{n}=\{(b_{1},\ldots,b_{n})\in L^{n}\mid\mathcal{D}\models\varphi(b_{1},\ldots,b_{n})\}.$
Once this has been established, we will obtain the desired statement, as the positive-existential ${\mathcal{L}_{\rm ring}}(k)$-definability in $L$ then follows via the usual quantifier-free interpretation of the finite-dimensional algebra $\mathcal{D}$ in $L$.
To establish the claim, consider $(b_{1},\ldots,b_{n})\in L^{n}$. Since $m/k$ is normal, $P(X)$ splits over $m$ (hence also over $M$) and any root of $P(X)$ generates $m$ over $k$ (hence also $M$ over $L$). It follows that $M$ is $L$-isomorphic to $L[X]/(Q_{i}(X))$ for each $i\in\{1,\ldots,g\}$. Since $\varphi$ is positive-existential, its validity is preserved under morphisms and direct products of rings, whereby we infer that
$(b_{1},\ldots,b_{n})\in S$ $\quad\Leftrightarrow\quad M\models\varphi(b_{1},\ldots,b_{n})$
$\quad\Leftrightarrow\quad\forall i\in\{1,\ldots,g\}:\frac{L[X]}{(Q_{i}(X))}\models\varphi(b_{1},\ldots,b_{n})$
$\quad\Leftrightarrow\quad\prod_{i=1}^{g}\frac{L[X]}{(Q_{i}(X))}\models\varphi(b_{1},\ldots,b_{n}).$
By the Chinese Remainder Theorem we have that $\mathcal{D}=\frac{L[X]}{(P(X))}\cong\prod_{i=1}^{g}\frac{L[X]}{(Q_{i}(X))}$ as $L$-algebras, from which the desired statement follows. ∎
@@ 6. Proof of the main theorem
Let $K_{0}$ denote the relative algebraic closure of $\mathbb{F}_{p}$ inside $F$. The hypotheses imply that $K_{0}$ is not algebraically closed.
@@
We use 6.1 to find $m\in\mathbb{N}$ such that $v(F(t^{1/p^{m}}))=v(F_{s}(t^{1/p^{m}}))$. Applying 2.1 with $k=l=\mathbb{F}_{p}(t)$, $L=F$ and $m=F(t^{1/p^{m}})$, we may replace $F$ by $F(t^{1/p^{m}})$ and $t$ by $t^{1/p^{m}}$ and thus assume from now on that $t$ is a uniformiser for the valuation $v$ on $F$.
Consider the residue field $Fv$ of $v$, which is a finite extension of $K$. Let $\ell$ denote the relative algebraic closure of $K_{0}$ in $Fv$; this is a finite extension of $K_{0}$. Since $K_{0}$ is an algebraic extension of $\mathbb{F}_{p}$ which is not algebraically closed, $K_{0}$ has a finite Galois extension of some prime degree $q$, whence the same holds for $\ell$. The field $F^{\prime}=F\otimes_{K_{0}}\ell$ is a function field in one variable over $\ell$, and the valuation $v$ on $F$ uniquely extends to a $\mathbb{Z}$-valuation on $F^{\prime}$ (which we also denote by $v$). We may again invoke 2.1 (this time with $k=\mathbb{F}_{p}(t)$, $l=K_{0}(t)$, $m=\ell(t)$, and $L=F$) to replace $F$ by $F^{\prime}$, and so we may assume without loss of generality that $v$ is a rational place, i.e. $Fv=K$. In particular, $F/K_{0}$ is a regular extension.
Let $f\in K_{0}[T]$ be an irreducible degree $q$ polynomial. Since any finite subextension of $K_{0}/\mathbb{F}_{p}$ is separable and hence generated by a single element, we may write $f(T)=g(\alpha,T)$ for some $g\in\mathbb{F}_{p}[A,T]$, $\alpha\in K_{0}$. Let $h(T)$ be the minimal polynomial of $\alpha$ over $\mathbb{F}_{p}$ and denote by $\alpha_{1},\ldots,\alpha_{k}$ the conjugates of $\alpha=\alpha_{1}$. The splitting field $G$ of $f$ over $F$ is identical to the splitting field of $g(\alpha_{i},T)$ for every $i\in\{1,\ldots,k\}$ and is a cyclic extension of $F$ of degree $l$; let $\sigma$ be a generator of its Galois group. We may consider the cyclic algebra $\mathcal{A}=(G,\sigma,t)$ which is by definition generated as an $F$-algebra by $G$ and an element $y$ subject to the relations $y^{q}=t$ and $xy=y\sigma(x)$ for all $x\in G$. In view of the previous discussion, by existentially quantifiying over the roots of $h(T)$ when describing $G$, we have that $\mathcal{A}$ is positive-existentially $\mathcal{L}$-interpretable in $F$ without parameters.

### v2

@@ 1. Introduction
This question, too, remains unresolved to this day. Anscombe and Fehm recently established a positive answer to a variant of 1.1, with $K(t)$ replaced by the field of formal Laurent series K((t))K(\!(t)\!) [AF16], whereas Tyrrell-Nic Dhonncha proved undecidability of a slightly more general decision problem in [TND25].
@@
Here, one may remark that undecidability of $\operatorname{Th}_{\exists}(\overline{\mathbb{F}_{p}}(t_{1},t_{2}),\mathbb{Z}[t_{1},t_{2}])$ was shown by Kim and Roush for $p>2$ [KR92], later generalized to include finite extensions of $\overline{\mathbb{F}_{p}}(t_{1},t_{2})$ as well by Eisenträger [Eis12]. Recently Becher, Dittmann and the author obtained more examples of function fields $F/K$ where $K$ contains $\overline{\mathbb{F}_{p}}$ and $\operatorname{Th}_{\exists}(F,R)$ is shown to be undecidable for some well-chosen ring of coefficients $R$ [BDD25], relying partially on the type of reduction techniques discussed in the present paper.
@@ 2. Preliminaries on definability
Consider an arbitrary field extension $L/K$ and a finite normal field extension $N/K$. Let $M$ be a common overfield of $L$ and $N$ over $K$ which is generated by $L\cup N$. If $S\subseteq M^{n}$ is positive-existentially ${\mathcal{L}_{\rm ring}}(K)$-definable over $M$, then $S\cap L^{n}$ is positive-existentially ${\mathcal{L}_{\rm ring}}(K)$-definable over $L$.
@@
Consider $\mathcal{D}=L\otimes_{K}N$, and denote by $\mathcal{D}^{\operatorname{\mathsf{red}}}$ its reduction (i.e. quotient modulo the nilradical). We observe the following properties:
  1. (1)
    Under the inclusion $N\to\mathcal{D}:x\mapsto 1\otimes x$, a $K$-basis of $N$ gets mapped to an $L$-basis of $\mathcal{D}$. In particular, $\mathcal{D}$ is a finite-dimensional $L$-algebra, and it has a basis whose structure constants all lie in $K$.
  2. (2)
    There exists $k\in\mathbb{N}$ such that every nilpotent element $a$ of $\mathcal{D}$ satisfies $a^{k}=0$.
  3. (3)
    There is a subset $G\subseteq\operatorname{Aut}(N/K)$ and a well-defined $L$-isomorphism $\Phi:\mathcal{D}^{\operatorname{\mathsf{red}}}\to M^{G}:$ such that $\Phi(a\otimes x)=(a\sigma(x))_{\sigma\in G}$ for all $a\in L$, $x\in N$.
The first property is immediate from standard properties of the tensor product. The second property follows because $\mathcal{D}$ is artinian and hence the nilradical is nilpotent (see e.g. [AM69, Proposition 8.4]).
For the third property: as $\mathcal{D}^{\operatorname{\mathsf{red}}}$ is a reduced artinian ring, it is isomorphic as an $L$-algebra to a finite product of field extensions of $L$, obtained by considering all quotients of $\mathcal{D}^{\operatorname{\mathsf{red}}}$ modulo maximal ideals. Clearly at least one of these quotients is equal to $M$, but since $N/K$ is normal, $\operatorname{Aut}(N/K)$ acts transitively on the spectrum of $\mathcal{D}^{\operatorname{\mathsf{red}}}$, from which the desired statement follows.
Now consider a positive-existential ${\mathcal{L}_{\rm ring}}(K)$-formula $\varphi(x_{1},\ldots,x_{n})$ such that $S=\{(b_{1},\ldots,b_{n})\in M^{n}\mid M\models\varphi(b_{1},\ldots,b_{n})\}$. We observe that
$S\cap L^{n}=\{(b_{1},\ldots,b_{n})\in L^{n}\mid\mathcal{D}^{\operatorname{\mathsf{red}}}\models\varphi(b_{1}\otimes 1,\ldots,b_{n}\otimes 1)\}.$
Indeed, in view of the $L$-isomorphism $\mathcal{D}^{\operatorname{\mathsf{red}}}\cong M^{G}$ from (3), there clearly exist $L$-homomorphisms $\mathcal{D}^{\operatorname{\mathsf{red}}}\to M$ and $M\to\mathcal{D}^{\operatorname{\mathsf{red}}}$, and validity of positive-existential formulas is preserved under homomorphisms.
The positive-existential ${\mathcal{L}_{\rm ring}}(K)$-definability of $S\cap L^{n}$ in $L$ then follows via the usual quantifier-free interpretation of the finite-dimensional algebra $\mathcal{D}$ (as ${\mathcal{L}_{\rm ring}}(K)$-structure) in $L$ (as ${\mathcal{L}_{\rm ring}}(K)$-structure), and the fact that there exists a natural number $k$ such that every nilpotent element $a\in\mathcal{D}$ satisfies $a^{k}=0$ (so for $a,b\in\mathcal{D}$ we have $\overline{a}=\overline{b}$ in $\mathcal{D}^{\operatorname{\mathsf{red}}}$ if and only if $(a-b)^{k}=0$ in $\mathcal{D}$). ∎
@@ 6. Proof of the main theorem
We use 6.1 to find $m\in\mathbb{N}$ such that $v(F(t^{1/p^{m}}))=v(F_{s}(t^{1/p^{m}}))$. Applying 2.1 we may replace $F$ by $F(t^{1/p^{m}})$ and $t$ by $t^{1/p^{m}}$ and thus assume from now on that $t$ is a uniformiser for the valuation $v$ on $F$.
Consider the residue field $Fv$ of $v$, which is a finite extension of $K$. In particular, the hypotheses on $K$ imply that $Fv$ does not contain an algebraically closed subfield. It follows that there exists some prime number $q$ and a finite extension $\ell/\mathbb{F}_{p}$ contained in $Fv$ such that $\ell$ has a finite Galois extension of degree $q$ not contained in $Fv$. The field $F^{\prime}=F\otimes_{F\cap\ell}\ell$ is a function field in one variable over $\ell$, and the valuation $v$ on $F$ uniquely extends to a $\mathbb{Z}$-valuation on $F^{\prime}$ (which we also denote by $v$). We may again invoke 2.1 to replace $F$ by $F^{\prime}$, and so we may assume without loss of generality that $\ell\subseteq F$.
Let $f\in\ell[T]$ be an irreducible degree $q$ polynomial. Since $\ell/\mathbb{F}_{p}$ is separable and hence generated by a single element, we may write $f(T)=g(\alpha,T)$ for some $g\in\mathbb{F}_{p}[A,T]$, $\alpha\in\ell$. Let $h(T)$ be the minimal polynomial of $\alpha$ over $\mathbb{F}_{p}$ and denote by $\alpha_{1},\ldots,\alpha_{k}$ the conjugates of $\alpha=\alpha_{1}$. The splitting field $G$ of $f$ over $F$ is identical to the splitting field of $g(\alpha_{i},T)$ for every $i\in\{1,\ldots,k\}$ and is a cyclic extension of $F$ of degree $l$; let $\sigma$ be a generator of its Galois group. We may consider the cyclic algebra $\mathcal{A}=(G,\sigma,t)$ which is by definition generated as an $F$-algebra by $G$ and an element $y$ subject to the relations $y^{q}=t$ and $xy=y\sigma(x)$ for all $x\in G$. In view of the previous discussion, by existentially quantifiying over the roots of $h(T)$ when describing $G$, we have that $\mathcal{A}$ is positive-existentially $\mathcal{L}$-interpretable in $F$ without parameters.